# New simple and accurate approximations for quintessence

### v2

@@ 2.5 DE Padé approximants
This is then inserted into (23a), which when used in (7) to obtain $f(N)$ leads to the same expression as (24) with (25) inserted. To obtain a better approximation for $\Omega_{\mathrm{DE}}(N)$ one can then go on to Padé approximants also involving $\beta$ in the expansion. [^6] These are given by $[1/2]_{\Omega_{\mathrm{DE}}}=\frac{T}{1+\gamma T-\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}$ and $[2/1]_{\Omega_{\mathrm{DE}}}=\frac{\gamma T+\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}{\gamma+\left[1+\frac{1}{2}(\gamma-1)\left(3+\beta+\gamma\right)\right]T}$. To express $T_{0}$ in terms of $\Omega_{\mathrm{DE}0}$ obtained in this manner requires solving quadratic equations in $T_{0}$, which leads to rather messy expressions. The approximation (24) turns out to be a fairly good approximation for quintessence. For this reason we will use the simple result in (25) for $T_{0}$ in (23b). This leads to that the expressions for $w_{\mathrm{DE}}$ in (23) can be written as follows: [^7] Note that one can use the series expansion in $T$ for $\Omega_{\mathrm{DE}}$ to obtain the following expression for $E_{\mathrm{DE}}^{2}$: $E_{\mathrm{DE}}^{2}\approx\Omega_{\mathrm{m}0}\exp(-3N)[1+T-(\gamma-1)T^{2}(1+{\textstyle{1\over 2}}(\gamma-1+\beta)T)+\dots]$ and use (25) for $T_{0}$. This expression can subsequently be used to obtain various Padé approximants to obtain good quintessence approximations for $E_{\mathrm{DE}}=E_{\phi}$ and then use $E_{\phi}(N)$ to compute, e.g., $w_{\mathrm{DE}}=w_{\phi}$, if one is so inclined.
$w_{\mathrm{DE}}$ $\approx[0/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right),$ (26a)
$w_{\mathrm{DE}}$ $\approx[1/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{\beta\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right),$ (26b)
@@
$w_{\mathrm{DE}0}$ $\approx w_{\infty}\left(\frac{1-\gamma\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}}\right),$ (27a)
$w_{\mathrm{DE}0}$ $\approx w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{1+(\beta-\gamma)\Omega_{\mathrm{DE}0}}\right).$ (27b)
Note that setting $\beta=\gamma-1$ in $[1/1]_{w_{\mathrm{DE}}}(T)$ yields $[0/1]_{w_{\mathrm{DE}}}(T)$ and that setting $\gamma=1$ in both $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ results in $w_{\mathrm{DE}}=w_{\infty}$, i.e., $\gamma=1$ leads to the $w$CDM model when $w_{\infty}\in(-1,0)$ and the $\Lambda$CDM model when $w_{\infty}=-1$.
@@
$f$ $=\frac{T}{T_{0}}\left(\frac{1+(\gamma-1)T_{0}}{1+(\gamma-1)T}\right)=\frac{\Omega_{\mathrm{m}0}e^{-3w_{\infty}N}}{1-\gamma\Omega_{\mathrm{DE}0}+(\gamma-1)\Omega_{\mathrm{DE}0}e^{-3w_{\infty}N}},$ (28a)
@@
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE},0}}{\Omega_{\mathrm{m}0}e^{3w_{\infty}N}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\mathrm{DE},0}},$ (29a)
$E_{\mathrm{DE}}^{2}$ $\approx\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\mathrm{DE}0}e^{-3(1+w_{\infty})N}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}}$ (29b)
@@
$\Omega_{\mathrm{DE}}$ $\approx\frac{\Omega_{\mathrm{DE},0}}{\gamma\Omega_{\mathrm{DE},0}+(1-\gamma\Omega_{\mathrm{DE},0})e^{3w_{\infty}N}},$ (30a)
@@
$E_{\mathrm{DE}}^{2}$ $\approx\Omega_{\mathrm{m}0}e^{-3N}\left(1+\frac{\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right).$ (30b)
It is trivial to express the above relations in the redshift $z$, which is done explicitly in section 4 where we investigate the accuracy of thawing and tracking quintessence approximations. To do so, we need $\gamma$ and $\beta$, which we derive in the next section, but before doing this let us briefly consider the CPL parametrization.
2.6 The Chevallier-Polarski-Linder parametrization
The so-called CPL parameters [25, 26] are often used in cosmology [^9] For a recent discussion of the CPL parametrization, see Wolf & Ferreira [33]. and parametrize $w_{\mathrm{DE}}$ according to
$w_{\mathrm{DE}}=w_{\mathrm{CPL}}=w_{0}+w_{a}\left(\frac{z}{1+z}\right)=w_{\infty}-w_{a}e^{N},$ (31)
where
$w_{\infty}=w_{0}+w_{a}.$ (32)
By using that $w_{a}=-w_{\mathrm{DE}}^{\prime}|_{N=0}$ we can easily compute the parameter $w_{a}$ from (26):
$w_{a}$ $=\frac{-3(\gamma-1)w_{\infty}^{2}\Omega_{\mathrm{DE}0}(1-\gamma\Omega_{\mathrm{DE}0})}{\Omega_{\mathrm{m}0}^{2}},$ (33a)
$w_{a}$ $=\frac{-3(\gamma-1)w_{\infty}^{2}\Omega_{\mathrm{DE}0}(1-\gamma\Omega_{\mathrm{DE}0})}{(1+(\beta-\gamma)\Omega_{\mathrm{DE}0})^{2}}.$ (33b)
Inserting (31) into (7) yields
$f=e^{-3w_{\infty}N}\cdot e^{-3w_{a}(1-e^{N})}=(1+z)^{3w_{\infty}}\cdot e^{-3w_{a}\left(\frac{z}{1+z}\right)},$ (34)
which results in that (13) and (12) yield
$\Omega_{\mathrm{CPL}}$ $=\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{DE}0}+\Omega_{\mathrm{m}0}e^{3w_{\infty}N}\cdot e^{3w_{a}(1-e^{N})}},$ (35a)
$E_{\mathrm{CPL}}^{2}$ $=\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\mathrm{DE}0}e^{-3(1+w_{\infty})N}\cdot e^{-3w_{a}(1-e^{N})},$ (35b)
and thereby
$\Omega_{\mathrm{CPL}}$ $=\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{DE}0}+\Omega_{\mathrm{m}0}(1+z)^{-3w_{\infty}}\cdot e^{3w_{a}\left(\frac{z}{1+z}\right)}},$ (36a)
$E_{\mathrm{CPL}}^{2}$ $=\Omega_{\mathrm{m}0}(1+z)^{3}+\Omega_{\mathrm{DE}0}(1+z)^{3(1+w_{\infty})}\cdot e^{-3w_{a}\left(\frac{z}{1+z}\right)}.$ (36b)
One can regard (31) as a series expansion in $e^{N}$, but this is inconsistent with the past matter dominant requirement $\Omega_{\mathrm{DE}}\rightarrow 0$ and that one of the CPL parameters can be interpreted as $w_{\infty}$, since this naturally leads to an expansion in $e^{-3w_{\infty}N}$, which is particularly pertinent as observations probe increasingly large redshift. CPL thus mixes this asymptotically motivated function (arising from the $w_{\infty}$ part in (31)) in $\Omega_{\mathrm{CPL}}$ and $E_{\mathrm{CPL}}^{2}$ with another function that comes from the $w_{a}$ part in (31). Due to this mixture of functions, especially the exponential part involving $\frac{z}{1+z}$, the CPL parametrization arguably does not lead to simpler expressions for observables than our $[0/1]_{w_{\mathrm{DE}}}(T)$ and $[1/1]_{w_{\mathrm{DE}}}(T)$ based results, and it certainly is not a good approximation for thawing and tracking quintessence.
3 Deriving the parameters $\gamma$ and $\beta$ for quintessence
@@ 3.1 Thawing quintessence
To obtain a dynamical system that is useful for deriving expressions for $\gamma$ and $\beta$ for thawing quintessence [^10] The concepts of thawing and freezing were defined by Caldwell and Linder (2005) [34] as follows: thawing is characterized by $w_{\varphi}\approx-1$ where $w_{\varphi}$ subsequently grows, i.e. $w_{\varphi}^{\prime}>0$, while $w_{\varphi}>-1$ and $w_{\varphi}^{\prime}<0$ holds for freezing. we introduce the following dimensionless bounded variable [^11] The variable $\Sigma_{\phi}$ was first introduced by Coley et al. (1997) [35] and Copeland et al. (1998) [36] whose $x$ is $\Sigma_{\varphi}$. Since then, $\Sigma_{\phi}$ (or $\phi^{\prime}$) is commonly used to describe scalar fields in cosmology, see, e.g., Urena-Lopez (2012) [37], equation (2.3), Tsujikawa (2013) [28], equation (16) and Alho and Uggla (2015) [38], equation (8). The reason for using the notation $\Sigma$ for the kernel is because $\Sigma_{\phi}$ plays a similar role as Hubble-normalized shear, which is typically denoted with the kernel $\Sigma$, see e.g. [39].
@@
and use $({\phi},\Sigma_{\phi},\Omega_{\phi})$ as the state vector together with the $e$-fold time $N$ [^12] The variables $u$ and $v$ in [30] yield slower convergence. To obtain a dynamical system for $({\phi},\Sigma_{\phi},\Omega_{\phi})$ we use that (14a) is equivalent to $\Omega_{\mathrm{DE}}^{\prime}=-3w_{\mathrm{DE}}\Omega_{\mathrm{DE}}(1-\Omega_{\mathrm{DE}})$, $\Omega_{\mathrm{DE}}=\Omega_{\phi}$, and that $w_{\mathrm{DE}}=w_{\phi}$ is given by
@@
Since thawing quintessence begins in the matter dominated regime, $\Omega_{\mathrm{m}}\approx 1$, when $w_{\varphi}\approx-1$ initially, it follows from (38) and from $\Omega_{\mathrm{m}}+\Omega_{\phi}=1$ that thawing quintessence begins when $\Sigma_{\varphi}^{2}\ll\Omega_{\phi}\ll\Omega_{\mathrm{m}}$. We then note that whenever $\lambda(\phi)$ is bounded, which always holds for some domain of $\phi$ for any scalar field potential $V(\phi)$, there is a line of matter dominated ‘Friedmann-Lemaître’ fixed points given by
@@
and hence $\Omega_{\mathrm{m}}=1$, parameterized by the constant values $\phi_{*}$. To determine what happens in the vicinity of $\mathrm{FL}^{\phi_{*}}$ we linearize the equations at $(\phi,\Sigma_{\phi},\Omega_{\phi})=(\phi_{*},0,0)$, which leads to
$\phi$ $\approx\phi_{*}+\frac{2\lambda_{*}}{9}Ce^{3N}-\frac{2}{3}Ke^{-\frac{3}{2}N},$ (42a)
$\sqrt{6}\Sigma_{\phi}$ $=\phi^{\prime}\approx\frac{2\lambda_{*}}{3}Ce^{3N}+Ke^{-\frac{3}{2}N}$ (42b)
@@
where $\lambda_{*}=\lambda(\phi_{*})$. It follows that thawing quintessence is associated with the unstable manifold of $\mathrm{FL}^{\phi_{*}}$, obtained by setting $K=0$. More precisely, thawing quintessence models consist of the 1-parameter set of unstable manifold orbits (i.e., solution trajectories) of $\mathrm{FL}^{\phi_{*}}$ and the open set of orbits that intermediately shadow these orbits; this open set of orbits consists of the orbits that before the thawing epoch are pushed toward $\mathrm{FL}^{\phi_{*}}$ by the stable manifold orbits of $\mathrm{FL}^{\phi_{*}}$ and subsequently toward the unstable manifold orbits of $\mathrm{FL}^{\phi_{*}}$, which therefore approximate these orbits during their thawing stage of evolution, for details, see [30] and [40].

### v3

@@ 2.5 DE Padé approximants and present initial data
This is then inserted into (26a), which when used in (7) to obtain $f(N)$ leads to the same expression as (28) with (29) inserted. At the prize of additional complexity, one can continue to produce Padé approximants for $\Omega_{\mathrm{DE}}$ that also involve the higher order term with $\beta$ in the expansion. [^6] The next two higher order Padé approximants, which involve $\beta$, are given by $[1/2]_{\Omega_{\mathrm{DE}}}(T)=\frac{T}{1+\gamma T-\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}$ and $[2/1]_{\Omega_{\mathrm{DE}}}(T)=\frac{\gamma T+\frac{1}{2}(\gamma-1)\left[1+(\beta-\gamma)\right]T^{2}}{\gamma+\left[1+\frac{1}{2}(\gamma-1)\left(3+\beta+\gamma\right)\right]T}$. This, however, leads to that expressing $T_{0}$ in terms of $\Omega_{\mathrm{DE}0}$ requires solving quadratic equations in $T_{0}$, which leads to rather messy expressions. The approximation (28) turns out to be a fairly good approximation for quintessence. For this reason we will use the simple result in (29) for $T_{0}$ in (26b). This leads to that the expressions for $w_{\mathrm{DE}}$ in (26) can be written as follows: [^7] Note that one can use the series expansion in $T$ for $\Omega_{\mathrm{DE}}$ to obtain the following expression for $E_{\mathrm{DE}}^{2}$: $E_{\mathrm{DE}}^{2}\approx\Omega_{\mathrm{m}0}\exp(-3N)[1+T-(\gamma-1)T^{2}(1+{\textstyle{1\over 2}}(\gamma-1+\beta)T)+\dots]$ and use (29) for $T_{0}$. This expression can subsequently be used to obtain various Padé approximants to obtain good quintessence approximations for $E_{\phi}(N)$ and then use $E_{\phi}(N)$ to compute, e.g., $w_{\phi}(N)$, if one is so inclined.
$w_{\mathrm{DE}}$ $=[0/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right),$ (30a)
$w_{\mathrm{DE}}$ $=[1/1]_{w_{\mathrm{DE}}}(T)=w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{\beta\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right),$ (30b)
@@
$w_{\mathrm{DE}0}$ $=w_{\infty}\left(\frac{1-\gamma\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}}\right),$ (31a)
$w_{\mathrm{DE}0}$ $=w_{\infty}\left(1-\frac{(\gamma-1)\Omega_{\mathrm{DE}0}}{1+(\beta-\gamma)\Omega_{\mathrm{DE}0}}\right).$ (31b)
@@
$\begin{split}f&=\frac{T}{T_{0}}\left(\frac{1+(\gamma-1)T_{0}}{1+(\gamma-1)T}\right)=\left(\frac{\Omega_{\mathrm{m}0}}{\Omega_{\mathrm{DE}0}}\right)\left(\frac{T}{1+(\gamma-1)T}\right)\\ &=\frac{\Omega_{\mathrm{m}0}e^{-3w_{\infty}N}}{1-\gamma\Omega_{\mathrm{DE}0}+(\gamma-1)\Omega_{\mathrm{DE}0}e^{-3w_{\infty}N}},\end{split}$ (32a)
@@
$\Omega_{\mathrm{DE}}$ $=\frac{\Omega_{\mathrm{DE}0}}{\Omega_{\mathrm{m}0}e^{3w_{\infty}N}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\mathrm{DE}0}},$ (33a)
$E_{\mathrm{DE}}^{2}$ $=\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\mathrm{DE}0}e^{-3(1+w_{\infty})N}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}}$ (33b)
@@
$\Omega_{\mathrm{DE}}$ $=\frac{T}{1+\gamma T}=\frac{\Omega_{\mathrm{DE}0}}{\gamma\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}},$ (34a)
@@
$\begin{split}E_{\mathrm{DE}}^{2}&=\Omega_{\mathrm{m}0}e^{-3N}\left(1+\frac{T}{1+(\gamma-1)T}\right)\\ &=\Omega_{\mathrm{m}0}e^{-3N}\left(1+\frac{\Omega_{\mathrm{DE}0}}{(\gamma-1)\Omega_{\mathrm{DE}0}+(1-\gamma\Omega_{\mathrm{DE}0})e^{3w_{\infty}N}}\right).\end{split}$ (34b)
Note the simple relationship between these expression and $\Lambda$CDM and $w$CDM, which are obtained by setting $\gamma=1$. Furthermore, eq. (34a) implies the following relations between the models in (34) and $\Lambda$CDM and $w$CDM at the present time: $\Omega_{\mathrm{DE}0}=\Omega_{\Lambda 0}/\gamma$ when $w_{\infty}=-1$ and $\Omega_{\mathrm{DE}0}=\Omega_{w0}/\gamma$ when $w_{\infty}\in(-1,0)$. As we shall see, thawing quintessence yields $\gamma>1$; on the other hand, depending on the scalar field potential, there exist tracking quintessence models with $\gamma>1$, $\gamma=1$ and $\gamma<1$.
Finally, note that it is trivial to express the above equations in the redshift $z$, since $z=\exp(-N)-1$. The above may be interpreted as exact cosmographic DE models, but this is not what we are interested in. Our objective is to derive approximations for solutions of the field equations for thawing and tracking quintessence. This requires deriving $\gamma$ and $\beta$ from the quintessence field equations, which, of course, involves the scalar field potential; this is then followed by inserting the results for $\gamma$ and $\beta$, expressed in terms of the scalar field potential and its derivatives, into the above DE formulas, which then are transformed from exact DE expressions to quintessence approximations.
3 Quintessence approximations
Before deriving expressions for the parameters $\gamma$ and $\beta$ for thawing and tracking quintessence, we begin with some background as regards quintessence. The concepts of thawing and freezing were defined by Caldwell and Linder (2005) [35] as follows: thawing is characterized by $w_{\varphi}\approx-1$ where $w_{\varphi}$ subsequently grows, i.e. $w_{\varphi}^{\prime}>0$, while $w_{\varphi}>-1$ and $w_{\varphi}^{\prime}<0$ holds for freezing. Some quintessence models go through several stages of thawing and freezing, see e.g. [36], and sometimes thawing quintessence and freezing quintessence therefore refer to a quintessence model at the present time, but this is not always the case. Quintessence is associated with that there exist particular quintessence solutions to the field equations that attract open sets of other quintessence solutions. Once a solution has been attracted to an ‘attractor’ quintessence solution it shadows it and thereby share its properties. It should be noted that these open sets of non-attractor quintessence solutions have a pre-quintessence stage during the matter dominated epoch before they are attracted to an attractor quintessence solution. During this stage of evolution the scalar field is effectively a test field and therefore not observationally relevant. To obtain an unambiguous nomenclature, we use the past asymptotic properties of the attractor solutions to denote the various types of quintessence.
For all scalar field potentials there is a 1-parameter set of thawing quintessence attractor solutions with $\phi=\phi_{*}=\mathrm{constant}$, $w_{\phi}=-1$, $\Omega_{\mathrm{m}}=1$, $\Omega_{\varphi}=0$ when $N\rightarrow-\infty$; for asymptotically inverse power law potentials with $V\propto\phi^{-p}$, $p>0$, when $\phi\rightarrow 0$, there is also a single tracker quintessence attractor solution with $-1<w_{\phi}=-2/(2+p)<0$, $\Omega_{\mathrm{m}}=1$, $\Omega_{\varphi}=0$ when $N\rightarrow-\infty$. For scalar field potentials that are sufficiently asymptotically exponentially steep there is a third type of quintessence: scaling freezing quintessence for which $w_{\phi}=0$ when $p_{\mathrm{m}}=0$ and $10<\Omega_{\mathrm{m}}<$ 1 when $N\rightarrow-\infty$. As a consequence this does not yield past continuous deformations of $\Lambda$CDM or $w$CDM and we will therefore only touch on this topic in appendix B. We thereby focus on thawing quintessence and tracking quintessence, which furthermore dominate what is observationally investigated in the literature. For more details and a rigorous global dynamical systems description of quintessence, see, e.g., [36], [37] and [38]. Next we derive dynamical systems that are adapted to obtaining the parameters $\gamma$ and $\beta$ for thawing and tracking quintessence, respectively.
@@ 3.1 Thawing quintessence
To obtain a dynamical system that is useful for deriving expressions for $\gamma$ and $\beta$ for thawing quintessence we introduce the following dimensionless bounded variable [^9] The variable $\Sigma_{\phi}$ was first introduced by Coley et al. (1997) [39] and Copeland et al. (1998) [40] whose $x$ is $\Sigma_{\phi}$. Since then, $\Sigma_{\phi}$ (or $\phi^{\prime}$) is commonly used to describe scalar fields in cosmology, see, e.g., Urena-Lopez (2012) [41], equation (2.3), Tsujikawa (2013) [30], equation (16) and Alho and Uggla (2015) [42], equation (8). The reason for using the notation $\Sigma$ for the kernel is because $\Sigma_{\phi}$ plays a similar role as Hubble-normalized shear, which is typically denoted with the kernel $\Sigma$, see e.g. [43].
@@
and use $({\phi},\Sigma_{\phi},\Omega_{\phi})$ as the state vector together with the $e$-fold time $N$ [^10] The variables $u$ and $v$ in [36] yield slower convergence. To obtain a dynamical system for $({\phi},\Sigma_{\phi},\Omega_{\phi})$ we use (20) with $\Omega_{\mathrm{DE}}=\Omega_{\phi}$ and that $w_{\mathrm{DE}}=w_{\phi}$ is given by
@@
Since thawing quintessence begins in the matter dominated regime, $\Omega_{\mathrm{m}}\approx 1$, when $w_{\varphi}\approx-1$ initially, it follows from (36) and from $\Omega_{\mathrm{m}}+\Omega_{\phi}=1$ that thawing quintessence begins when $\Sigma_{\phi}^{2}\ll\Omega_{\phi}\ll\Omega_{\mathrm{m}}$. We then note that whenever $\lambda(\phi)$ is bounded, which always holds for some domain(s) of $\phi$ for any scalar field potential $V(\phi)$, there is a line of matter dominated ‘Friedmann-Lemaître’ fixed points in the present dynamical systems formulation given by
@@
and hence $\Omega_{\mathrm{m}}=1$, parametrized by the constant values $\phi_{*}$. To determine what happens in the vicinity of $\mathrm{FL}^{\phi_{*}}$ we linearize the equations at $(\phi,\Sigma_{\phi},\Omega_{\phi})=(\phi_{*},0,0)$, which leads to
$\phi$ $\approx\phi_{*}+\frac{2\lambda_{*}}{9}Ce^{3N}-\frac{2}{3}De^{-\frac{3}{2}N},$ (40a)
$\sqrt{6}\Sigma_{\phi}$ $=\phi^{\prime}\approx\frac{2\lambda_{*}}{3}Ce^{3N}+De^{-\frac{3}{2}N}$ (40b)
@@
where $\lambda_{*}=\lambda(\phi_{*})$. It follows that thawing quintessence is associated with the unstable manifold of $\mathrm{FL}^{\phi_{*}}$, obtained by setting $D=0$. More precisely, thawing quintessence models consist of the 1-parameter set of unstable manifold orbits (i.e., solution trajectories) of $\mathrm{FL}^{\phi_{*}}$ and the open set of orbits that intermediately shadows these orbits; this open set of orbits consists of the orbits that before the thawing epoch are pushed toward $\mathrm{FL}^{\phi_{*}}$ by the stable manifold orbits of $\mathrm{FL}^{\phi_{*}}$ and subsequently toward the unstable manifold orbits of $\mathrm{FL}^{\phi_{*}}$, which therefore approximate these orbits during their thawing stage of evolution, for details, see [36] and [38].
@@
Inserting the result for $\gamma$ in (44) into (30a) and (34) yields our lowest order thawing quintessence approximations for $w_{\phi}$, $\Omega_{\phi}$ and $E_{\phi}$, while inserting the results for $\gamma$ in (44) and $\beta$ in (45) into (30b) and (33) gives our corresponding most accurate approximations. Note that these formulas are valid for any scalar field potential and any $\phi_{*}$, but the approximations are only accurate for observationally viable solutions, i.e., models for which $\epsilon(\phi_{*})$ and $\eta(\phi_{*})$ are typically ${\cal O}(1)$ or less. Based on section 2 where we now replace the subscript DE with ϕ we obtain the following approximations for thawing quintessence: Throughout we use eq. (29) to connect with initial data, i.e.
$T_{0}=\frac{\Omega_{\phi 0}}{1-\gamma\Omega_{\phi 0}},$ (46)
where $\Omega_{\phi 0}$ is observationally determined, which leads to that there is no error at all for our $\Omega_{\phi}$ approximations at the present time, [^11] A given value of $\Omega_{\phi 0}$ describes a surface in the state space $(\phi,\Sigma_{\phi},\Omega_{\phi})$. This surface determines where $N=0$ for the various orbits in the state space. although there will be small errors for both $w_{\phi 0}$ and $E_{\phi 0}$ from the $w_{\phi}\approx[1/1]_{\mathrm{DE}}(T)$ based approximation, which we will focus on. On the other hand, computing $w_{\phi 0}$ numerically for one of the solutions from the 1-parameter set of thawing quintessence attractor solutions for a specific scalar field potential and using eq. (27a) for $T_{0}$ results, of course, in no deviation at all from the numerically calculated value for $w_{\phi 0}$.
Using that $w_{\infty}=-1$ for thawing quintessence leads to $T=T_{0}\exp(3N)$ where we recall that $T_{0}=\frac{\Omega_{\phi 0}}{1-\gamma\Omega_{\phi 0}}$ (see e.q. (46)), while (30b) and (33) yield our master equations for thawing quintessence:
$w_{\phi}$ $\approx[1/1]_{w_{\mathrm{DE}}}(T)=-\left(1-\frac{(\gamma-1)T}{1+\beta T}\right),$ (47a)
$\Omega_{\phi}$ $\approx\frac{\Omega_{\phi 0}}{\Omega_{\mathrm{m}0}e^{-3N}\left(\frac{1+\beta T}{1+\beta T_{0}}\right)^{\frac{\gamma-1}{\beta}}+\Omega_{\phi 0}},$ (47b)
$E_{\phi}^{2}$ $\approx\Omega_{\mathrm{m}0}e^{-3N}+\Omega_{\phi 0}\left(\frac{1+\beta T_{0}}{1+\beta T}\right)^{\frac{\gamma-1}{\beta}},$ (47c)
where we recall that
$\gamma$ $=1+\left(\frac{2}{3}\right)^{3}\epsilon_{*},$ (47d)
$\beta$ $=\frac{3}{5}-\left(\frac{2}{3}\right)^{3}\epsilon_{*}+\frac{4}{15}\eta_{*},$ (47e)
where the slow-roll parameters $\epsilon_{*}$ and $\eta_{*}$ are given by $\epsilon_{*}=\epsilon(\phi_{*})=\left.\frac{1}{2}\left(V_{,\phi}/V\right)^{2}\right|_{\phi=\phi_{*}}$ and $\eta_{*}=\eta(\phi_{*})=\left.V_{,\phi\phi}/V\right|_{\phi=\phi_{*}}$; the 1-parameter set of values $\phi_{*}$ are the freezing values of $\phi$ during the matter dominated regime where $\Omega_{\phi}\approx 0$, where each value yields a thawing quintessence attractor solution, where these solutions collectively describe an open set of solutions that shadow them during the epoch when quintessence affects the spacetime structure. The simpler, but less accurate approximations, although with similar accuracy as previous approximations in the literature, based on $w_{\phi}\approx[0/1]_{w_{\mathrm{DE}}}(T)$, are obtained by setting $\beta=\gamma-1$ in the above expressions. Recall also, that setting $\gamma=1$, which corresponds to that $\epsilon_{*}=0$, yields the $\Lambda$CDM solution. To express our approximations in the redshift $z$, just use the simple relation $z=\exp(-N)-1$. The application of the above master equations is illustrated with a specific example, a quintessential $\alpha$-attractor inflation potential, in section 5.